(b) If the record of time starts when x = +A and v = 0, determine expressions for the displacement, velocity, and acceleration of the oscillating mass along the x-axis at any time t later. (Your expression should be in terms of the variable t and other numerical values. Assume any numerical values in your expression are in standard SI units, but do not enter units into your expression.) x =? Acos(?t) Incorrect: Your answer is incorrect. How is the x component of the location of the particle on the reference circle related to the displacement of the oscillating mass on the x-axis? v =? ?A?sin(?t) Incorrect: a =? ?A?2cos(?t) Incorrect: Your answer is incorrect. (b) If the record of time starts when x = +A and v
Answer #1

a) let at time t, angle made by the position vector with +x axis w*t

so, cos(w*t) = x/A

==> x = A*cos(w*t)

b) v = dx/dt

= A*(-sin(w*t)*w)

= -A*w*sin(w*t)

c) a = dv/dt

= -A*w*(cos(w*t)*w)

= -A*w^2*cos(w*t)

These answers must be according to the give conditions. at t = 0, x = +A

please check your initial conditions. for the given initial condtions are these equations are correct.